dudley and Dolores collect hubcaps. Dudley needs 18 to match Dolores' collection. Together they have 184. How many does each have now?
Found 2 solutions by JulietG, lwsshak3:
Answer by JulietG(1812)

Y + 18 = D [Dudley needs 18 to match Dolores' collection.]
Y + D = 184
Substitute the known value of D from the first equation into the second.
Y + (Y+18) = 184
2Y + 18 = 184
Subtract 18 from each side.
2Y = 166
Y = 83. Dudley has 83.
.
Let's make certain:
83 + 18 =101 (Dolores' amount)
101 (Dolores) + 83 (Dudley) = 184
Success!

Answer by lwsshak3(11628)

dudley and Dolores collect hubcaps. Dudley needs 18 to match Dolores' collection. Together they have 184. How many does each have now?
***
let x=Dolores' collection
x-18=Dudley's collection
x+x-18=184
2x=202
x=101
x-18=83
Dolores' collection=101 hubcaps
Dudley's collection=83 hubcaps